OpenStudy (anonymous):

you have 3/6 cup of oatmeal in your cupboard . a recipe calls for 1/4 cup of oatmeal .How much oatmeal would you have left if you made the cookies ?

OpenStudy (anonymous):

\[\frac{3}{6}-\frac{1}{4}\] \[\frac{2}{4}-\frac{1}{4}\] \[\frac{1}{4}\]